# Is Smaller Always Faster? Tradeoffs in Compressing Self-Supervised Speech Transformers

### v3

@@ II Related Work
Recent studies have proposed task-agnostic compression methods that aim to preserve the general usability of speech representations. DistilHuBERT [8] distills a 12-layer HuBERT into a shallow 2-layer student by aligning hidden representations using cosine and $\ell_{1}$ losses. FitHuBERT [9] retains the model depth but reduces the width of attention heads, feedforward layers, and convolutional channels to achieve a thinner architecture. LightHuBERT [10] introduces a two-stage training process: first distilling a large model into a supernet, then applying once-for-all training to support a range of subnetwork configurations. ARMHuBERT [11] and STaRHuBERT [12] share a similar architecture with FitHuBERT but achieve superior performance through masking distillation and temporal relation distillation, respectively. Beyond distillation-based approaches, several task-agnostic, pruning-based methods have been proposed for speech self-supervised models [22, 23]. Additionally, Lin et al. [24] propose a method that dynamically determines whether to perform early exiting during feature extraction.
@@ IV Design of Experiments
To understand compression tradeoffs, we pre-train and compress 12-layer Transformers using the entire LibriSpeech dataset, which contains 960 hours of read speech. We evaluate representations on two downstream tasks: phoneme recognition using the 100-hour subset of LibriSpeech, and speaker identification using VoxCeleb1. Following SUPERB [32], pre-trained models are frozen, and downstream classifiers learn a weighted sum of all layers. All pruning experiments use pre-training loss as a proxy for downstream performance, and we show loss curves and key observations to motivate our design choices.
@@ IV-A MelHuBERT
We train MelHuBERT-10ms and MelHuBERT-20ms with masked prediction [4, 33] for two pre-training stages. For the first stage, we first run k-means with 512 clusters on log Mel features extracted from the 960-hour subset of LibriSpeech. We then train a 12-layer Transformer to predict the cluster labels of each frame for MelHuBERT-10ms or to predict every other frame for MelHuBERT-20ms. For the second stage, we run k-means with 512 clusters on the ninth layer hidden representations of the first stage model. We predict each frame for both MelHuBERT-10ms and MelHuBERT-20ms in the second stage. We choose a masking strategy based on [25], using a 7% masking probability with a mask of 10 frames (100 ms) for MelHuBERT-10ms and a 14% masking probability with a mask of 5 frames (100 ms) for MelHuBERT-20ms. For the first stage, both MelHuBERT-10ms and MelHuBERT-20ms are trained on 8 V100 GPUs with Adam for 200 epochs. For the second stage, we train MelHuBERT-10ms for 200 epochs and MelHuBERT-20ms for 100 epochs. We use an effective batch size of 32 and a learning rate of $10^{-4}$ for both stages. Dropout of 10% is applied after the multiplication of matrices, such as query, key, and value matrices, and after FC1 and FC2.
@@ IV-B Weight Pruning
For weight pruning, we iteratively prune individual weights based on their $\ell_{1}$ for all weights and biases in linear layers of Transformers. We keep track the exponential moving average of the loss with a decay of 0.9998, and if the loss does not change much (within 0.001) compared to the one 15,000 steps before, pruning is triggered. We use a pruning schedule that is aggressive when the network is dense, and mild when the network is sparse. In particular, we prune 20% until 80% density, 10% until 50% density, 5% until 35% density, 2.5% until 30% density, 1% until 10% density, and 0.5% until 5% density. We use a batch size of 4 and a learning rate of $10^{-5}$. The pruning schedule follows [28, 34] with minor modifications.
@@ IV-C Head Pruning
For head pruning, we use a 25% of the training data to compute the scores of each head. Similarly, since the $\ell_{1}$ norm varies across layers, as opposed to pruning a fixed number for each layer, we normalize the scores of heads within each layer and prune the heads by comparing them altogether. As for all iterative pruning approaches, we train the model for a fixed 25,000 steps after pruning, with learning rate $10^{-5}$ and a batch size of 4.
@@ IV-E Knowledge Distillation
For knowledge distillation, we explore a 2-layer and a 6-layer student network. Following prior works [10, 11], we study several design decisions, including changing the masking strategy, initializing with MelHuBERT layers, and the temperature when training with KL divergence. Based on our study, neither applying masking on the input nor initializing with MelHuBERT layers lead to improvement. Hence, we simply minimize KL divergence as our distillation loss without masking and initializing.
@@ V-A Downstream performance
For knowledge distillation, we find that both the 2-layer and 6-layer students are able to retain the performance of teachers on speaker identification, with the 6-layer ones performing even better than the teachers. Both students are not able to match the teachers’ performance on phone recognition, with the 2-layer one lacking significantly behind. The worse performance seems to suggest that knowledge distillation might not be an effective compression approach. However, as we will see in later sections, once we put all the results in context, knowledge distillation actually performs favorably compared to others.
@@ V-B MACs, parameters, and runtime
While Figure 1 and Figure 2 shows how well each approach maintains performance after compression, it does not show the benefit of compression. Density is a relative measure, making it difficult to compare across approaches. To evaluate the benefit of compression across approaches, we decide to use three metrics, the number of multiply-accumulation operations (MACs) per one second speech, the number of parameters, and the runtime measured on a single GPU. The three metrics are related. Reduction in MACs represents the theoretical speed-up of inference when parallel computation is not available. In the presence of a GPU, the reduction in MACs does not always translate to faster runtime, as it depends on the structure of the pruned networks. Reduction in model parameters usually leads to a reduction in MACs and can improve generalization. As we have seen in weight pruning and head pruning (Figure 1), there is a period along the pruning process where the pre-trained loss and downstream performance both improve. Results of all approaches evaluated on the three metrics are shown in Figure 3. We only focus on phone recognition, but it shows a similar trend for speaker identification.
@@
Knowledge distillation (with a 2-layer and a 6-layer student) achieves a significant reduction on all three metrics. In fact, knowledge distillation provides the best tradeoff among all approaches. We attribute this to the the number of GPU kernel calls. Each GPU kernel call has a significant overhead, and computing with fewer layers not only reduces MACs per one second speech, but also saves the overhead. For example, a model pruned with low-rank approximation can have a similar MACs per one second speech compared to a 6-layer student, but the runtime of the pruned model is significantly slower than the 6-layer student.
Finally, inspired by the results of knowledge distillation, we provide a baseline that only takes the first two layers and the first six layers of the pre-trained Transformers. This amounts to stopping the forward process at the second and the sixth layer. This approach involves no additional training and is arguably the simplest baseline. Combined with other approaches in Figure 3, we now have a better picture what the worst and the best are able to achieve.
@@ V-D Combining compression techniques
The results of combining compression techniques are shown in Figure 4. The combined approach is strictly better than knowledge distillation alone, and is close to weight pruning on MACs per one secone speech and numbers of parameters. In addition, the resulting models are also strictly better than DistilHuBERT and FitHuBERT. Compared to dedicated approaches, simple combination of compression techniques is able to produce strong results.
@@ VI Conclusion
Based on the weight pruning results, we can conclude that there exists small but strong self-supervised models that can produce representations that generalize across tasks. Weight pruning, even the vanilla version used in this work, is in fact able to find models much smaller than the recent approaches. However, smaller models do not necessarily run faster. Even though runtime is correlated with MACs, the characteristics of the hardware (such as the amount of parallelization) plays a much important role. In other words, if the goal is to improve runtime, reducing the number of GPU kernel calls is much more effective than reducing the size of the kernel calls. We hope that this study encourages future work to consider multiple dimensions when designing and evaluating compression strategies.

### v4

@@ II Related Work
Recent studies have proposed task-agnostic compression methods that aim to preserve the general usability of speech representations. DistilHuBERT [8] distills a 12-layer HuBERT into a shallow 2-layer student by aligning hidden representations using cosine and $\ell_{1}$ losses. FitHuBERT [9] retains the model depth but reduces the width of attention heads, feedforward layers, and convolutional channels to achieve a thinner architecture. LightHuBERT [10] introduces a two-stage training process: first distilling a large model into a supernet, then applying once-for-all training to support a range of subnetwork configurations. ARMHuBERT [11] and STaRHuBERT [12] share a similar architecture with FitHuBERT but achieve superior performance through masking distillation and temporal relation distillation, respectively. Beyond distillation-based approaches, several task-agnostic, pruning-based methods have been proposed for speech self-supervised models [22, 23, 24]. Additionally, Lin et al. [25] propose a method that dynamically determines whether to perform early exiting during feature extraction.
@@ IV Design of Experiments
To examine compression tradeoffs, we pre-train and compress 12-layer Transformers on the full 960-hour LibriSpeech dataset. Representations are evaluated on phoneme recognition (LibriSpeech 100-hour subset) and speaker identification (VoxCeleb1), following SUPERB [33] with frozen encoders and weighted-sum layer aggregation. Pre-training loss serves as a proxy for downstream performance in pruning experiments.
@@ IV-A MelHuBERT
We adopt MelHuBERT-10ms and -20ms [26], pre-trained in two stages with masked prediction [4, 34]. Stage 1 uses k-means (512 clusters) on log Mel features to train a 12-layer Transformer for frame-wise prediction (10ms) or every-other-frame prediction (20ms). Stage 2 applies k-means on the 9th-layer features from Stage 1, with frame-wise prediction in both variants.
[…]
@@ IV-B Weight Pruning
For weight pruning, we iteratively prune individual weights based on their $\ell_{1}$ for all weights and biases in linear layers of Transformers. We keep track of the exponential moving average of the loss with a decay of 0.9998, and if the loss does not change much (within 0.001) compared to the one 15,000 steps before, pruning is triggered. We use a pruning schedule that is aggressive when the network is dense, and mild when the network is sparse. In particular, we prune 20% until 80% density, 10% until 50% density, 5% until 35% density, 2.5% until 30% density, 1% until 10% density, and 0.5% until 5% density. We use a batch size of 4 and a learning rate of $10^{-5}$. The pruning schedule follows [29, 35] with minor modifications.
@@ IV-C Head Pruning
For head pruning, we use 25% of the training data to compute the scores of each head. Similarly, since the $\ell_{1}$ norm varies across layers, as opposed to pruning a fixed number from each layer, we normalize the scores of heads within each layer and prune the heads by comparing them all together. As for all iterative pruning approaches, we train the model for a fixed 25,000 steps after pruning, with learning rate $10^{-5}$ and a batch size of 4.
@@ IV-E Knowledge Distillation
For knowledge distillation, we explore a 2-layer and a 6-layer student network. Following prior works [10, 11], we study several design decisions, including changing the masking strategy, initializing with MelHuBERT layers, and the temperature when training with KL divergence. Based on our study, neither applying masking on the input nor initializing with MelHuBERT layers leads to improvement. Hence, we simply minimize KL divergence as our distillation loss without masking and initializing.
@@ V-A Downstream performance
For knowledge distillation, we find that both the 2-layer and 6-layer students are able to retain the performance of teachers on speaker identification, with the 6-layer ones performing even better than the teachers. Neither student is able to match the teachers’ performance on phone recognition, with the 2-layer one lagging significantly behind. The worse performance seems to suggest that knowledge distillation might not be an effective compression approach. However, as we will see in later sections, once we put all the results in context, knowledge distillation actually performs favorably compared to others.
@@ V-B MACs, parameters, and runtime
While Figure 1 and Figure 2 show how well each approach maintains performance after compression, they do not show the benefit of compression. Density is a relative measure, making it difficult to compare across approaches. To evaluate the benefit of compression across approaches, we decide to use three metrics, the number of multiply-accumulation operations (MACs) per one second speech, the number of parameters, and the runtime measured on a single GPU. The three metrics are related. Reduction in MACs represents the theoretical speed-up of inference when parallel computation is not available. In the presence of a GPU, the reduction in MACs does not always translate to faster runtime, as it depends on the structure of the pruned networks. Reduction in model parameters usually leads to a reduction in MACs and can improve generalization. As we have seen in weight pruning and head pruning (Figure 1), there is a period along the pruning process where the pre-training loss and downstream performance both improve. Results of all approaches evaluated on the three metrics are shown in Figure 3. We only focus on phone recognition, but it shows a similar trend for speaker identification.
@@
Knowledge distillation (with a 2-layer and a 6-layer student) achieves a significant reduction on all three metrics. In fact, knowledge distillation provides the best tradeoff among all approaches. We attribute this to the number of GPU kernel calls. Each GPU kernel call has a significant overhead, and computing with fewer layers not only reduces MACs per one second speech, but also saves the overhead. For example, a model pruned with low-rank approximation can have a similar MACs per one second speech compared to a 6-layer student, but the runtime of the pruned model is significantly slower than the 6-layer student.
Finally, inspired by the results of knowledge distillation, we provide a baseline that only takes the first two layers and the first six layers of the pre-trained Transformers. This amounts to stopping the forward process at the second and the sixth layer. This approach involves no additional training and is arguably the simplest baseline. Combined with other approaches in Figure 3, we now have a better picture of what the worst and the best are able to achieve.
@@ V-D Combining compression techniques
The results of combining compression techniques are shown in Figure 4. The combined approach is strictly better than knowledge distillation alone, and is close to weight pruning on MACs per one second speech and numbers of parameters. In addition, the resulting models are also strictly better than DistilHuBERT and FitHuBERT. Compared to dedicated approaches, simple combination of compression techniques is able to produce strong results.
@@ VI Conclusion
Based on the weight pruning results, we can conclude that there exist small but strong self-supervised models that can produce representations that generalize across tasks. Weight pruning, even the vanilla version used in this work, is in fact able to find models much smaller than the recent approaches. However, smaller models do not necessarily run faster. Even though runtime is correlated with MACs, the characteristics of the hardware (such as the amount of parallelization) play a much more important role. In other words, if the goal is to improve runtime, reducing the number of GPU kernel calls is much more effective than reducing the size of the kernel calls. We hope that this study encourages future work to consider multiple dimensions when designing and evaluating compression strategies.